core: 0 of 1 pt HW Score: 75%, 4.5 of 6 p Bus Econ 5.4.50 Question Help A n w employee charged S 7640 on his credit card to relocate for his rst job. After noticing that the interest rate for his balance was 18% compounded monthly, he stopped charging on that account. He wishes to pay off his balance in 2 years using automatic payments sent at the end of each month. a. What monthly payment must he make to pay off the account at the end of 2 years? b. How much total interest will he have paid? a. What monthly payment must he make to pay off the account at the end of 2 years? (Round to the nearest cent as needed)
Answer #1

(a). The formula used to calculate the fixed monthly payment (P) required to fully amortize Outstanding dues of $ of L over a term of n months at a monthly interest rate of r is P = L[r(1 + r)n]/[(1 + r)n - 1].

Here, L = 7640 , n = 2*12 = 24 and r = 18/100 = 0.18 so that P = 7640*0.18*(1.18)24/[(1.18)24 -1] = 1375.20(53.10900627)/ 52.10900627 = 1401.59 ( on rounding off to the nearest cent).

Thus, the person should make a monthly payment of $ 1401.59 to pay off the account at the end of 2 years.

(b). The total payment made by the person is 24* $ 1401.59 = $ 33638.16 so that the interest paid by him is $ 33638.16 -$ 7640 = $ 25998.16.